OpenStudy (anonymous):

Richard surveyed his class and found that 14 out of 26 students read the sports section first.write the ratio as a fraction in the simplest form and as a decimal to the nearest thousandth.

OpenStudy (anonymous):

14/26 = 7/13 = around 0.538